MATH 41 Probability& Stalsles Homewors Fat calorics 190 220 Sat, fat (819LD 0 a) Construct a scatter plot for the dat b) Compute the correlation coefficien for the data c) Describe the lincar relationship betem Fat calorlies and Sat fa d) Find the equation of the ngion line the d
Answer #1

(a)

30 gm 25 gm 20 gm 15 gm 10 gm 500 550 400 450 5 gm 300 350 200 250 150 Fat Calories Series 4 Series 5 Series 6 . Series Series 2 Series 3 - (best line) - (best line) -(best line) -(best line) - (best line) -(best line) meta-chart.com

(b)

X Values
∑ = 2040
Mean = 340
∑(X - Mx)2 = SSx = 96600

Y Values
∑ = 97
Mean = 16.167
∑(Y - My)2 = SSy = 292.833

X and Y Combined
N = 6
∑(X - Mx)(Y - My) = 5280

R Calculation
r = ∑((X - My)(Y - Mx)) / √((SSx)(SSy))

r = 5280 / √((96600)(292.833)) = 0.9927

(c) This is a strong positive correlation, which means that high fat calories go with high saturated fat(and vice versa).

(d)

Sum of X = 2040
Sum of Y = 97
Mean X = 340
Mean Y = 16.1667
Sum of squares (SSX) = 96600
Sum of products (SP) = 5280

Regression Equation = ŷ = bX + a

b = SP/SSX = 5280/96600 = 0.05466

a = MY - bMX = 16.17 - (0.05*340) = -2.41718

ŷ = 0.05466X - 2.41718